# On the Oscillations in Cournot Games with Best Response Strategies

### v1

@@ 3 Periodic Oscillation
When the number of firms with consistently positive production during the two-period stable oscillation is $k_{1}$. By distinguishing the value of $k_{1}$, three types of equilibrium can be achieved:
@@
    When $k_{1}=0$, in one of the two periods, the production of each firm is zero. In this case, the existence of the solution can be verified in $O(n)$ time and the unique solution can be obtained in $O(n)$ time.
@@
    The production of each of $k_{1}<3$ firms remains consistently positive, while in one of the two periods, the production of each remaining firm is zero. The existence of such an oscillation can be verified in $O(n)$ time. There could in a infinite number of oscillations in this case and a solution in this case can be found in $O(n)$ time.
@@
    Oscillations occur in the case of three firms. The existence of such an oscillation can be verified in $O(1)$ time. There could in a infinite number of oscillations in this case and a solution in this case can be found in $O(1)$ time.

### v2

@@ 3 Periodic Oscillation
Let $k_{1}$ denote the number of firms with consistently positive production during a two-period stable oscillation. Depending on the value of $k_{1}$, the system exhibits three distinct equilibrium types:
@@
    Case $k_{1}=0$: In one of the two periods, all firms have zero production.
    • Equilibrium existence can be verified in $O(n)$ time.
    • The unique equilibrium can be computed in $O(n)$ time.
@@
    Case $30<k_{1}<$ 3: Exactly $k_{1}$ firms maintain positive production in both periods, while all others have zero production in one period.
    • Oscillation existence can be verified in $O(n)$ time.
    • Infinitely many oscillations may exist.
    • A solution can be find in $O(n)$ time.
@@
    Case $k_{1}\geq 3$ : Stable oscillations emerge among three or more firms.
    • Oscillation existence can be verified in $O(1)$ time.
    • Infinitely many oscillations may exist.
    • A solution can be find in $O(1)$ time.